# A3: Android Agent Arena for Mobile GUI Agents

### v1

@@ 4 Experiments
Test Subset Test Level Succ. Rate
IDD High 69.6
Low 92.1
Category High 51.8
Unseen Low 84.4
App High 56.8
Unseen Low 83.0
Task High 73.7
Unseen Low 88.5
Table 4: Static frame evaluation results on AndroidControl test set.
4.1 Finetuned Agent
We train an agent based on InternVL2-8B (Chen et al., 2024) using both the AMEX and AndroidControl datasets. The agent is then tested on the AndroidControl static frame test set as well as the A3 dynamic arena. Table4 presents the evaluation results on the AndroidControl test set, which is divided into four subsets: IDD, category unseen, app unseen, and task unseen. Each static frame evaluation is further categorized into two levels. The high-level evaluation provides only an overall task goal, such as “Open the profile tab in YouTube,” while the low-level evaluation includes additional step-wise instructions for the current screen, such as “Click the profile icon at the bottom right of the screen.” Naturally, low-level tasks are easier than high-level ones, as they offer more specific guidance.
Although the agent performs relatively well on the AndroidControl static test set, we observe poor performance in real-world scenarios, as shown in Table 5 and Table 6. Upon analyzing unsuccessful executions in A3, we identify that the history of actions might be a significant factor. (i) Previous studies have demonstrated that action history improves performance in static frame evaluations, as it is derived from ground-truth annotations and provided as input to the agent, which informs subsequent processes. However, in A3, real-world scenarios lack ground-truth action histories, which is a huge domain gap between training data and real-world evaluation. If the agent relies on its own action history, a single mistake can confuse the agent and cascade into a series of errors. (ii) In static evaluations, errors in action predictions do not affect subsequent states or the ground-truth history. However, in real-world evaluation, a wrong action can lead to an incorrect state, which might provide misleading information and requires the agent to self-correct, either by going back or starting over. Unfortunately, our trained agent currently lacks this self-correction capability.
Another challenge is the impracticality of information query tasks. This issue arises because no existing dataset provides this type of data during collection, leading the agent to fail in answering any queries.
Agent Test Level Succ. Rate
InternVL2 Easy 23.4%
Medium 5.6%
Hard 2.0%
GPT-4o Easy 9.9%
Medium 1.4%
Hard 0.0%
AppAgent Easy 30.8%
Medium 7.0%
Hard 2.0%
Table 5: Dynamic evaluation results on A3 by difficulty level.
4.2 Business-level LLM
Agent Test Category Succ. Rate
InternVL2 Operation 17.1%
Single-frame Query 0.0%
Multi-frame Query 0.0%
GPT-4o Operation 5.7%
Single-frame Query 2.0%
Multi-frame Query 0.0%
AppAgent Operation 20.0%
Single-frame Query 6.0%
Multi-frame Query 0.0%
Table 6: Dynamic evaluation results on A3 by task category.
We also employ the business-level LLM GPT-4o as an agent in A3. However, as highlighted in the previous Set-of-Mark (SoM) study (Yang et al., 2023), GPT-4o performs poorly when tasked with directly outputting or inputting coordinates. To address this, we apply the SoM technique to help identify element positions and coordinates. While SoM improves GPT-4o’s performance on CLICK actions, it continues to struggle with other actions, such as SCROLL. Despite these limitations, GPT-4o demonstrates the ability to complete some information query tasks due to its inherent capabilities, even without fine-tuning.
We also evaluate the performance of AppAgent (Zhang et al., 2023a), which includes an exploration phase before task automation and articulates its reasoning and planning during execution. These features notably enhance its effectiveness as an agent. AppAgent not only significantly outperforms the original GPT-4o but also surpasses our finetuned agent, likely due to the advanced capabilities of the GPT model underlying AppAgent.
4.3 Error Cases
@@
  • Perform CLICK at wrong coordinate. This the most common error case where if the coordinate is wrong, either the screen doesn’t change (click at nothing), or it goes to a wrong state, which requires self-correct ability to get back on the track. This type of errors is mainly due to the insensibility of LLMs to digits.
  • Perform meaningless action. This happens when an agent doesn’t know what to do on the current state. For instance, it generates WAIT when the state is fully loaded and requires a CLICK or SCROLL. This is possibly due to the weak planning capability of the agent.
@@
  • Can’t stop. We also see some cases where the agent already finishes the task but it still performs more actions that ruin the process.
5 Limitations
@@
  • Business-level LLMs such as GPT-4o and Gemini can only evaluate whether the whole task is complete, but cannot judge whether the sub-goals are completely in the action chain. This remains to the future work to discover a better way to autonomously evaluate the agent performance other than success rate.
6 Discussion
@@ Appendix A Appendix
A.1 Task Examples
@@
A.2 Prompts
The simplified prompt for GPT-4o to generate evaluation function code (Section 3.4.2) is as followed:
Given a task and a xml file of the screen, generate an evaluation code for the task to check whether the task is completed correctly. Think of the conditions carefully, especailly different elements and whether to use regex expression.
The task is ’{task}’. And here is the xml of the screen: {xml}.
The evaluation function should start with the following code:
‘‘‘python
The evaluation function should return True if the task is completed correctly and False otherwise.
A base xml parser is also provided for you to use:
‘‘‘python
for example you can use the following code to evaluate the task: Search Michael Jackson on YouTube.’
‘‘‘python
Output the generated code with only code, not wrapped in a markdown.
Listing 1: Prompt for evaluation function code generation.
@@ A.2 Prompts
Given a task and a xml file of the screen, judge whether the task is completed correctly.
The task is ’{task}’. And here is the xml of the screen: {xml}.
To correctly judge the performance, you should consider element states such as selected and activated to check whether specific condition is satisfied. You also need to consider the content-desc attributes of elements to check the result correctness.
Answer with "yes" or "no".
Listing 2: Prompt for evaluation function code generation.
A.3 Finetuned Agent
We use the 8B model of InternVL2 as the base model. We train the model on 8 A100 GPU for 27 hours on AndroidControl and AMEX, following the default finetuning settings.
A.4 Demonstrations

### v2

@@ 4 Experiments
4.1 Evaluation
We evaluate six models or agents and compare their performance using the A3 evaluation system. Three of these models (InternVL2 (Chen et al., 2024), Qwen-2.5-VL (Team, 2025), UI-TARS (Qin et al., 2025)) are fine-tuned or pre-trained specifically for GUI tasks, while the other three (GPT-4o-2024-11-20, Claude-3.5-sonnet, AppAgent (Zhang et al., 2023a)) are commercial LLMs or models that leverage commercial LLMs.
We select the 7B versions of Qwen2.5-VL and UI-TARS, along with the 8B version of InternVL2, to ensure a comparable agent size. Qwen2.5-VL is a generalist Multimodal Large Language Model (MLLM) pre-trained on GUI-related data, while UI-TARS further fine-tunes Qwen2-VL on GUI-related data with enhancements for reasoning and screen understanding. Both agents are evaluated following their official inference guidelines. Additionally, we apply supervised fine-tuning to InternVL2 on AMEX  (Chai et al., 2024) and AndroidControl (Li et al., 2024) datasets to serve as a baseline agent (further fine-tuning details are provided in Appendix A.5). For GPT-4o, we incorporate Set-of-Mark (SOM) (Yang et al., 2023) to improve element recognition and selection, while for Claude, we adapt a specialized prompt modified from the “Claude-computer-use” prompt to better suit mobile scenarios. AppAgent is used directly without any modifications.
Table 4 presents the main evaluation results for the six agents. “Func SR” refers to the task success rate as evaluated by predefined functions, “LLM SR” represents the task success rate as assessed by commercial LLMs, and “ESAR” stands for the essential states achieved rate.
The table reveals that agents pre-trained or fine-tuned on GUI tasks generally outperform generalist models, suggesting that commercial generalist models lack sufficient adaptation for GUI-specific tasks. Even though Claude’s “computer use” model demonstrates significantly better performance than GPT-4o, it still faces a substantial domain gap between desktop environments and mobile scenarios. AppAgent, on the other hand, leverages GPT-4o’s basic image understanding capabilities and incorporates additional exploration and indexing techniques, resulting in a significant performance improvement compared to the original GPT-4o and Claude’s “computer use” model. Moreover, UI-TARS achieves considerably higher success rates than Qwen2.5-VL, underscoring the value of task-specific fine-tuning and the impact of reasoning data. Finally, InternVL2, as a baseline fine-tuned agent, demonstrates that even with minimal fine-tuning, smaller models can outperform well-designed commercial LLMs on GUI-specific tasks.
We observe that commercial LLM evaluations consistently yield higher success rates (SRs) compared to function-based evaluations. Upon analyzing their evaluation processes, we find that LLMs apply looser criteria than functions which often enforce strict conditions, such as requiring a specific final state or a predefined sequence of actions. In contrast, commercial LLM evaluations align more closely with human judgment, as they assess tasks based on essential states and the overall sequence of actions. Additionally, the ESAR metric offers a more detailed evaluation, even when the task success rate is all 0.0 for difficult tasks. ESAR reflects an agent’s ability to break down complex tasks into simpler essential states, with higher ESAR values indicating better performance. This trend aligns with the patterns observed in task SRs.
We further validate the function-based and commercial LLM evaluation results through human observation. The average accuracy of commercial LLM evaluations is approximately 93%, closely matching the estimate calculated in Section 3.4.2. This high level of accuracy, confirmed by human validation, strongly supports the reliability of commercial LLM evaluations in real-world scenarios. Additionally, the observation that LLM evaluations are occasionally more accurate than function-based evaluations can be attributed to the reasons outlined in Section 3.4.1.
4.2 Error Case Analysis
@@
  • Perform CLICK at wrong coordinate. If the coordinate is wrong, either the screen doesn’t change (click at nothing), or it goes to a wrong state, which requires self-correct ability to get back on the track. This type of errors is mainly due to the insensibility of LLMs to digits.
  • Wrong action planning. This happens when an agent predicts completely wrong action type. With reasoning and thinking, UI-TARS still suffers from the action prediction when a task is complex.
@@
  • “Stop” issue. Some agents such as Qwen2.5-VL often stops too early and other agents such as UI-TARS fail to stop even if the task is finished successfully.
5 Discussion
@@ Appendix A Appendix
For anonymization, the url is temporarily at https://anonymous.4open.science/r/AVD-Control-Evaluation-F758.
A.1 App Selection
We select the apps from different categories in Google Play Store, where only two categories (News and Shopping) contains 2 apps. Those apps satisfy two conditions: (i) can be installed and run normally on any emulators; (ii) in top 15 list of the category. The first condition is to ensure the apps can run on any devices, because many apps don’t support x86 platforms. The second condition ensures the popularity of the apps to represent the daily scenario.
A.2 Task Generation and Examples
The tasks are all created by system designers and function coder. Each coder is in charge of an app and creates its most common usages and tasks based on the app functions. For example, tasks of shopping apps will contain essential states such as “search for products”, “add to wishlist”, “add to cart”, “sort by price”, etc. The difficulty of tasks can be controlled by different number of essential states. For example, “sort by price” or “filter to under \$100” can be optional to modify the task difficulty level. Coders are also required to keep the ratio of three levels at around 5:3:2. Then the coder will determine the level of the task by executing a “Gound Truth” trace and saving all the information such as screenshot, XML and action history. These information are used to design the evaluation functions. The functions are then validated by three correct traces and three wrong traces.
The following are task examples:
@@ A.2 Task Generation and Examples
A.3 Essential States
The essential states of a task can be generated by two approaches: (i) human definition and (ii) LLM generation (GPT-4o-2024-11-20). To ensure the correctness and precision, users can choose the human definition. For example, if the task is “Open Booking.com and search for one-way flight from Beijing to Shanghai on Feb 20. Select two adults.” The user can define the essential states: ["Device is in Booking.com app", "One-way flight is selected", "Departure city is set to Beijing", "Arrival city is set to Shanghai", "Departure date is set to Feb 20", "Number of adults is set to 2", "Search results are displayed"]. To autonomously execute the evaluation process, users can choose LLM generation. We also validate the LLM generated essential states of 30 tasks, 10 for each difficulty level. The validation result is that the correctness of LLM generated essential states reaches 100%, which means it can perfectly split a task into essential states as humans do. The prompt is as followed:
[…]
Listing 1: Prompt for essential states generation.
A.4 Prompts
The simplified prompt for GPT-4o to evaluate essential states (Section 3.4.2) is as followed:
Given a list of UI elements, an image consisting of screenshots
and a list of states, judge whether one or more states are correctly
[…]
You are required to reply with the states that are correctly achieved
without any changes. If none of the states are correctly achieved,
[…]
If the "answer" is correct, reply with the the "answer" attribute "yes" and
[…]
Also provide the reason for your answer.
Listing 2: Prompt for essential states evaluation.
@@ A.4 Prompts
[…]
Answer with only "yes" or "no". The reason should consider the conflicting elements or wrong states
Listing 3: Prompt for final state evaluation.
A.5 Finetuned Agent
We use the 8B model of InternVL2 as the base model. We train the model on 8 A100 GPU for 27 hours on AndroidControl and AMEX, following the default finetuning settings. The data is directly used for supervised finetuning, such as “Given the task: xxxxx, your action history: xxxxx, predict next action on the screen <image>” and “CLICK<coor>100, 100</coor>”.
A.6 Demonstrations